OpenStudy (anonymous):

solve for x 2(x-4)-(x+2)=(x+2)(x-4)

OpenStudy (anonymous):

2x-8-x-2=x^2-2x-8 x^2-x-2=0 and factor

OpenStudy (anonymous):

huh i dont understand where you got the 0 and factor from

OpenStudy (anonymous):

you need to solve both sides of the equation, it's simple algebra.. add/subtract like terms. 2x-8-x-2=x^2-2x-8 +8 +8 2x -x-2=x^2-2x +2x +2x -x^2+3x-2=0 (I skipped a step but i moved everything to the left side of the equation.) -(x-1)(x-2)=0 x=1 x=2